An angry construction worker throws his wrench downward from a height of 128 feet with an initial velocity of 32 feet per seconds is givin S(t) = -16t^2 - 32t + 128.
A)what is the height of the wrench after 1 second?
B) H ow long does it take for the wrench to reach the ground?

Answer by stanbon(75887)

An angry construction worker throws his wrench downward from a height of 128 feet with an initial velocity of 32 feet per seconds is givin
S(t) = -16t^2 - 32t + 128.
A)what is the height of the wrench after 1 second?
S(1) = -16-32+128 = 80 ft.
---------------------------------
B) How long does it take for the wrench to reach the ground?
The height is zero when the wrench is on the ground.
Solve: -16t^2 - 32t + 128 = 0
Divide thru by -16 to get:
t^2+2t-8 = 0
(t+4)(t-2) = 0
Positive solution:
t = 2 seconds
=============================
Cheers,
Stan H.